the dosage recommendation for medication A is 5.03 mg A/Kg of body weight. How many cc of medication should be given to a patient who weighs 144 lbs if the solution contains 18.7 mg of medication per cc.
Answer #1

Given,

dosage recommendation for medication A =5.03 mg/kg of body weight

Body weight of patient =144lbs

Since ,

1lbs =0.4536 kg

then,

144lbs =0.4536*144=65.31 kg

Now,

since 1 kg of body weight requires 5.03mg of A

then,65.31 kg of body weight requires 5.03*65.31 mg of A

=328.50mg

Now,

the solution contains 18.7mg per cc (means 18.7mg in 1cc)

then, for 1mg =1/18.7 cc

for 328.50 mg =(1/18.7)*328.50 cc

=17.56 cc

So, 17.56 cc of medication is given to the patient of 144lbs .